OpenStudy (katielong):

@phi another:

OpenStudy (katielong):

\[r=7\sqrt{t}\] would be: \[t=\frac{ r ^{2} }{ 7 }\] ?

OpenStudy (phi):

I would divide both sides by 7 (to get the t by itself) \[ \frac{r}{7}= \sqrt{t} \] now square both sides (to get rid of the square root) what do you get?

OpenStudy (katielong):

t=r^2/7 would both the top and bottom be squared or just the top?

OpenStudy (phi):

square means multiply by itself: \[ \frac{r}{7}\cdot \frac{r}{7}= \sqrt{t} \cdot \sqrt{t}\]

OpenStudy (katielong):

so 2r on top?

OpenStudy (phi):

When multiplying fractions, you multiply top times top, and bottom times bottom r*r can be written as \( r^2\) It is short-hand, and it's more useful for r*r*r*r*r, written as \(r^5\) notice that r+r = 2r (1 r plus another r gives you 2 r's)

OpenStudy (katielong):

ahhhh sooo.. r^2/49?

OpenStudy (phi):

yes